OpenStudy (anonymous):

About how much pressure do the feet of a 6000-kg elephant exert on the ground? Assume each foot has an area of 0.10 square meters. Show your work.

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

What have you done on your own ?

OpenStudy (anonymous):

@jess.white

OpenStudy (anonymous):

I thought it was this but it was wrong. Pressure is force/area therefore Force:6000-kg Area is 4*(.1m^2)= .4m^2 force/area = 6000-kg/.4m^2=15000kg/m^2 @ShailKumar

OpenStudy (anonymous):

Good but you have put force = 6000 Kg but that is mass ! So, if you are given mass, any idea about force ?

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

OK, force = weight = mass times g. Take g = 10 m/s^2. And then do the calculations again as you did earlier.

OpenStudy (anonymous):

could you help please :(

OpenStudy (anonymous):

OK. Pressure = 6000*10/0.4 N/m^2. Now calculate.

OpenStudy (anonymous):

@jess.white Could you get the result ?

OpenStudy (anonymous):

I dont know how to get it

OpenStudy (anonymous):

6000*10/0.4 = ?

OpenStudy (anonymous):

is it 2,400?

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

No, There is division. It's 150000 N/m^2

OpenStudy (anonymous):

oh ok i see now that is what i just did

OpenStudy (anonymous):

can you tell me the whole equation so i can write it for the show my work part

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

P = F/ A = mg/A

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok i am abut to post last question

OpenStudy (anonymous):

Sure...